OpenStudy (anonymous):

1.|x| + 5 = 18 5 or –5 13 or –13 18 or –18 23 or –23 2. |x + 3| < 5 –8 < x < 2 –2 < x < 8 3 < x < 5 –3 < x < 5 3.|–3n| – 2 = 4 2 or –2 3 or –3 4 or –4 6 or –6 4. |f| – 2/3 5/6 5/6 or -5/6 2/3 or -2/3 1 1/2 or -1 1/2 no solution Please help? I suck at Equations and Inequalities

OpenStudy (ash2326):

@Jessica_7897 Do you know how absolute value function is defined?

OpenStudy (anonymous):

@ashish9277 I believe I know what absolute value is

OpenStudy (ash2326):

ok \[|x|+5=18\] so \[\pm x+5=18\] Can you solve this from here?

OpenStudy (anonymous):

x equals 13 or -13?

OpenStudy (ash2326):

yes you're correct :D

OpenStudy (ash2326):

Next is \[|x+3|<5\] Can you try this?

OpenStudy (anonymous):

Is the answer –3 < x < 5?

OpenStudy (anonymous):

1.x+5=18 x=13 2.x+3<5 x<3 3. 3n-2=4 3n=6 n=2 4.if it is |f| – 2/3 =5/6 then f-2/3=5/6 f=2/3+5/6 f=4/6+5/6 f=11/6 f=1 5/6 hope it helps

OpenStudy (anonymous):

1;13 or –13 2;-8<x<2 3;=-2,2 4;no solution

OpenStudy (anonymous):

@robert2 Thank you so much!